# Toward a Physical Theory of Intelligence

### v1

@@ 5.2 Conjecture: Intelligence as Constrained Entropy-Production Maximisation
We now advance a conjecture that situates intelligent organisation within a broader physical principle governing open systems.
In open systems supplied with persistent free-energy gradients, dynamical structures that increase the achievable rate of entropy production tend to be selected, subject to dynamical, structural, and resource constraints. When dissipation pathways are unconstrained, entropy production may be maximised through unstructured or turbulent dynamics. However, when constraints limit access to gradients, actuation bandwidth, or the rate of irreversible information processing, maximising entropy production favours the formation of internal predictive structure and control. Under such constraints, intelligence emerges as a mechanism for selectively unlocking and routing environmental free energy through available pathways, thereby increasing entropy export while minimising the internal irreversible cost required to do so.
This conjecture reframes intelligence not as a violation of the second law of thermodynamics, but as one of its most refined consequences. Intelligent systems do not suppress entropy production; they organise it. By shaping internal dynamics to preserve structure and deploying dissipation only where it produces maximal environmental effect, intelligent agents expand the set of accessible dissipation pathways and increase the rate at which environmental free energy is released under constraints.
5.3 Entropy, Information, and the Arrow of Intelligence
The metriplectic formulation together with the preceding conjecture supports viewing intelligence as a process of constrained entropy maximisation. In open systems supplied with persistent free-energy gradients, irreversible information processing enables an agent to selectively access and exploit those gradients, converting them into goal-directed work that ultimately manifests as entropy production in the environment. Internal informational structure is maintained not in opposition to entropy production, but precisely because it allows entropy to be produced more effectively under constraints, by reducing wasted dissipation and concentrating irreversible interventions along goal-aligned degrees of freedom.
From this perspective, goal-directed work is not an extraneous objective superimposed on physical dynamics, but the mechanism by which entropy production is organised. An intelligent system routes entropy production through structured pathways that are aligned with its goals, thereby maximising the achievable rate of entropy export subject to limitations on dissipation pathways, actuation bandwidth, and the physical cost of irreversible information processing. Intelligence therefore increases when a larger fraction of the total entropy production enabled by the system is realised through controlled, task-relevant interactions with the environment rather than through unstructured or redundant dissipation.
@@ 5.3 Entropy, Information, and the Arrow of Intelligence
This view situates intelligent behaviour within a broader tendency observed in nonequilibrium physics. Many open systems evolve toward dynamical configurations that increase entropy production subject to their constraints, as documented across physical and biological settings [10, 23]. Intelligent agents are no exception. What distinguishes them is not a departure from this tendency, but the manner in which it is realised: by developing internal predictive structure and control, they expand the set of accessible dissipation pathways and increase the rate at which environmental free energy can be released in a structured, goal-aligned way.
@@
An instructive analogy comes from applied mathematics and classical physics. A projectile follows a parabolic trajectory not because the system “chooses” that path, but because it is the unique solution consistent with the governing equations and constraints. In the same sense, intelligent behaviour arises as the natural resolution of the reversible-dissipative constraints governing an open system interacting with its environment. Given persistent free-energy gradients and limits on how dissipation can be effected, the formation of internal informational structure and goal-directed control is not incidental but inevitable: it is the most effective way for the system to maximise entropy production under the constraints it faces.
Under this interpretation, intelligence is not an anomaly within physical law, but a direct and highly structured expression of the second law in complex, open systems. Because irreversible information processing defines a direction of influence that cannot be undone [18, 23], intelligence is inherently time-asymmetric. It exists only in systems whose dynamics distinguish past from future through dissipative structure, and whose internal organisation allows entropy production to be increasingly concentrated into goal-aligned, environmentally mediated work. In this sense, the emergence and progression of intelligent behaviour reflects the same physical drive that underlies nonequilibrium processes themselves: the tendency of constrained systems to realise ever more effective pathways for entropy production.
@@ 6 Consciousness and the Epistemic Limits of Intelligent Systems
In the preceding sections, we introduced a physically grounded definition of intelligence as the amount of goal-directed work produced per nat of irreversibly processed information. An intelligent system must therefore transform information in a manner that minimises irreversible loss while maximising useful work. This asymmetry imposes a fundamental constraint: to sustain high intelligence over a given horizon, a system must preserve a non-trivial fraction of its internal informational structure. Preserved information reduces cumulative irreversibility by allowing the system to reuse, rather than continually destroy, informational distinctions that remain relevant to future behaviour.
Persistent internal structure may take many forms. Reflexive behaviours rely on long-lived mappings such as stable sensorimotor relationships, while more sophisticated agents preserve richer encodings, including memories, latent structure, internal constraints, and regularities of the environment. Such preserved structure forms the substrate of self-modelling: the process by which an agent maintains information about its own dynamics and its interaction with the environment across time. In this sense, the emergence of self-modelling reflects the retention of information that continues to constrain future behaviour rather than being overwritten through irreversible operations. Related views of consciousness as arising from persistent internal models or self-referential structure have been explored in theoretical biology and cognitive science, though without a physical efficiency-based formalisation [32].
To formalise this distinction, we introduce a quantity dual to intelligence, which we term consciousness. Consciousness is defined as the amount of goal-directed work enabled per nat of preserved information over a finite horizon $T$. This measure quantifies how effectively preserved information constrains future behaviour under the dynamics, complementing the intelligence metric, which quantifies the efficiency of irreversible information processing. Together, these quantities characterise the twin informational requirements of adaptive systems: the transformation of new information and the preservation of structure necessary for long-horizon control.
The remainder of this section develops the consequences of this view. We formalise self-modelling as the internal process by which preserved information constrains future behaviour and introduce the corresponding consciousness metric. We then show that, within this framework, increasing intelligence over extended horizons correlates with increasing consciousness. Finally, we establish that self-modelling systems possess unavoidable epistemic limitations: because preserved information is necessarily a coarse-grained encoding of internal dynamics, no physically realisable agent can form a complete model of itself. This situates self-modelling and its epistemic boundaries within the same physical framework governing information, conserved quantities, and intelligent behaviour.
6.1 Persistent Information and Consciousness
The definition of intelligence introduced in §3.3 characterises an agent by the amount of goal-directed work it produces per nat of irreversibly processed information. To complement this quantity, we now examine informational structure that is preserved across the evolution of the agent’s internal dynamics. When such persistent information constrains future behaviour in a way that supports long-horizon control, it functions as a form of self-modelling. Under the present framework, the preservation of information that improves goal-directed behaviour forms the basis for a physically grounded notion of consciousness.
The internal state $X_{t}$ evolves according to the coupled agent-environment dynamics of §2 through the boundary variables $(\Gamma_{X}(X_{t}),\Gamma_{E}(E_{t}))$. Because all interaction with the environment is mediated through these ports, any information that remains available to the agent across time must encode structural relationships relevant to future behaviour under the agent’s own dynamics and control. Persistent information therefore reflects features of the agent-environment coupling that remain invariant over the behavioural horizon and continue to constrain future trajectories.
Under the CCE assumptions of §2.3.1, each encoding $\ell\in L$ is realised by a metastable equivalence class of physical trajectories, represented by a basin or invariant set $B_{\ell}\subset P$ under a control configuration $c\in C$, as specified by the encoding map $f:L\times C\to\mathcal{P}(P)$. Such equivalence classes need not correspond to fixed-point attractors; they may equally represent limit cycles, oscillatory phases, or other invariant dynamical structures. Information is preserved over time precisely when the system remains within the same equivalence class and is destroyed only when an irreversible merge collapses previously distinguishable classes.
In this setting, physical configurations identify encodings up to basin-scale fluctuations: there exists a generally partial, operational readout map $d:P\times C\to L$ that is well defined whenever the system’s state lies within a metastable equivalence class, in the sense that
$p\in B_{\ell}(c)\;\Rightarrow\;d(p,c)=\ell\quad\text{(up to fluctuations within the class).}$ (60)
We therefore write the instantaneous encoding as
$\ell_{t}:=d(P_{t},C_{t})\in L.$ (61)
An informational distinction is said to be preserved over a horizon $[t,t+T]$ if the system’s internal trajectory remains within the same metastable equivalence class under admissible control throughout the interval, i.e. if no irreversible merge of encoding classes occurs. Preserved information may be expressed continuously or only become operationally accessible over time (e.g. through oscillatory phase or other dynamical modes), but remains recoverable without additional irreversible computation.
Let $\mathscr{I}^{\mathrm{pres}}_{t}(T)\subseteq L$ denote the set of encodings whose equivalence classes remain unmerged over the horizon. The total preserved information is then defined as
$I_{\mathrm{preserved}}(T):=H\big(\mathscr{I}^{\mathrm{pres}}_{t}(T)\big),$ (62)
measured in nats.
We now introduce a quantity dual to intelligence, which we term consciousness. Over a finite horizon $T$, define
$\kappa_{T}:=\frac{W_{\mathrm{goal},T}}{I_{\mathrm{preserved}}(T)}.$ (63)
Consciousness $\kappa_{T}$ therefore quantifies how effectively preserved information constrains future behaviour so as to enable goal-directed work. Intuitively, reserved structure that is irrelevant to increasing goal-directed work does not increase $\kappa_{T}$ but reduces it.
This definition identifies consciousness as a physically measurable property of an embodied agent. Intelligence characterises the efficiency with which irreversibly processed information is converted into useful work, whereas consciousness characterises the efficiency with which preserved information supports the same process. Because persistence reduces the need for fresh irreversible computation, consciousness and intelligence arise from complementary aspects of the agent’s information dynamics.
@@ 6.2 Emergence of Consciousness in Intelligent Systems
Intelligence increases when an agent produces goal-directed work while reducing the amount of irreversible information processing required to do so. In environments containing reusable temporal or structural regularities, an agent that repeatedly reconstructs the same distinctions incurs unnecessary irreversible cost. Under the GENERIC dynamics of §2.4, such redundant computation manifests as excessive flow through the dissipative channel $R\nabla\Xi$, increasing both generalized entropy production and irreversible-information rate. Conversely, an agent that invests resources into forming persistent informational structure can reuse those distinctions across time, reducing the need for repeated dissipative intervention and thereby increasing intelligence.
Recall that the reversible component $J\nabla H$ preserves internal informational structure, while the dissipative component $R\nabla\Xi$ implements irreversible interventions that break time-reversal symmetry. The instantaneous rate of goal-directed work decomposes as
$\dot{W}_{\mathrm{goal}}=-\,\langle\nabla G,\,J\nabla H\rangle+\langle\nabla G,\,R\nabla\Xi\rangle-\partial_{t}G.$ (64)
Both components may contribute to useful work through interaction with the environment; however, work obtained via dissipative intervention necessarily incurs irreversible information loss. For a fixed goal-directed work requirement, intelligence therefore increases by reducing unnecessary reliance on dissipative computation rather than by eliminating dissipation altogether.
Proposition 7.1 formalizes this principle by showing that, among all metriplectic dynamics achieving a prescribed mean rate of goal-directed work, the long-run irreversible-information rate is minimised when dissipative computation is confined to task-critical interventions and internal reversible dynamics are used to preserve and reuse informational structure. This result does not imply that all useful work must be carried by the reversible flow; rather, it identifies the selective concentration of irreversibility as the mechanism by which intelligence is increased.
The formation of persistent informational structure is therefore the mechanism through which an agent reduces redundant irreversible computation while maintaining or increasing useful work. Building such structure typically requires transient investment of resources, but once established it allows future behaviour to be guided by reversible internal dynamics without repeated dissipative cost. Redundant recomputation through $R\nabla\Xi$ is avoided, irreversible-information flow is reduced, and intelligence increases.
Under the definition introduced in §6.1, this persistent, behaviourally relevant informational structure constitutes the agent’s consciousness. The metriplectic framework thus reveals consciousness not as an optional add-on, but as a necessary consequence of increasing intelligence over extended horizons: any agent that improves its efficiency of irreversible information use in a structured environment must accumulate and maintain internal information that supports reuse, prediction, and long-horizon control. In this sense, consciousness is the persistent informational organisation required to sustain intelligent behaviour under physical constraints.
6.3 Epistemic Limits of Consciousness
Intelligent systems operating in structured environments must preserve internal informational structure in order to sustain high intelligence over extended horizons. Under the framework developed in §6.1, this preserved structure constitutes the system’s consciousness and increases as intelligence increases. We now show that finite preserved-information capacity necessarily implies epistemic incompleteness: no physically realizable system can form a complete model of its own internal dynamics.
Under the CCE framework, preserved information corresponds to membership in metastable equivalence classes of internal trajectories. Let $L^{\mathrm{pres}}_{t}(T)\subseteq L$ denote the set of encoding classes that remain unmerged over the horizon $[t,t+T]$. Because preserved information is finite,
$|L^{\mathrm{pres}}_{t}(T)|<\infty,$ (65)
and distinct internal configurations or trajectories must therefore map to the same preserved encoding class.
Proposition 6.1 (State incompleteness).
If the preserved encoding set $L^{\mathrm{pres}}_{t}(T)$ is finite, then there exist internal state distinctions that cannot be represented within the system’s preserved information.
Proof sketch.
Because $L^{\mathrm{pres}}_{t}(T)$ is finite, there exist distinct internal states or trajectories that belong to the same preserved equivalence class. Any representation based solely on preserved information must identify these distinct states, and therefore cannot encode all true distinctions about the system’s internal configuration. ∎
Thus, there are always physically real distinctions about the system’s internal state that lie beyond the resolution of its preserved informational structure.
The same limitation applies to the system’s predictions about its own future evolution. Let $K_{X}$ denote the internal transition kernel, and for each internal state $X$ define $F_{k}(X):=\mathsf{Law}(X_{t+k}\mid X_{t}=X)$. If the dynamics are non-degenerate—i.e. there exist $X_{1}\neq X_{2}$ with $F_{k}(X_{1})\neq F_{k}(X_{2})$ for some kk—but both states belong to the same preserved equivalence class, then no representation based solely on preserved information can recover the correct future distributions for both states.
Proposition 6.2 (Dynamical incompleteness).
@@ 6.3 Epistemic Limits of Consciousness
If preserved information is finite and the internal dynamics are non-degenerate, then there exist distinctions about the system’s future evolution that cannot be inferred from its preserved informational structure.
These results establish a physical form of epistemic incompleteness. Just as no sufficiently expressive formal system can encode all truths about itself, no physically realizable system with finite preserved information can encode all true distinctions about its own internal dynamics. This incompleteness arises not from logical paradox, but from the irreversibility and coarse-graining inherent in physical information processing.
Because preserved information is necessarily limited by the system’s finite capacity for maintaining metastable distinctions, the expressive power of consciousness is bounded. Systems with greater preserved information capacity may represent more than weaker systems, but all remain subject to the same structural limitation: complete self-models are physically unattainable.
@@ 7 Intelligence and the Dynamics of the Brain
Two empirical observations are particularly relevant. First, neural oscillations are ubiquitous: coherent rhythms appear across frequency bands in EEG, LFP, MEG, and intracellular recordings, coordinating population activity across spatial and temporal scales [7, 12, 30]. In our framework, oscillatory dynamics implement relatively low-dissipation information transformations by routing activity along structured manifolds of the state space. Their phase-coherent coordination reduces the irreversible component of computation, thereby lowering the irreversible-information rate required to sustain a given level of goal-directed work.
@@
Second, neural dynamics exhibit signatures of near-criticality: power-law statistics, neuronal avalanches, and long-range temporal correlations suggest that cortical systems operate near the edge of a dynamical critical regime [4, 26]. Such regimes maximise sensitivity, dynamic range, and information propagation while maintaining coherence—properties that align naturally with our efficiency-based metrics relating entropy flow, work extraction, and computational flexibility.
@@
In this section we argue that oscillatory and near-critical dynamics jointly position the brain in an informationally efficient operating regime under the proposed framework, illustrating how biological neural systems realise the physical principles underlying intelligent behaviour.
7.1 Oscillatory Dynamics
A central implication of the framework introduced in §2 is that intelligent systems benefit from internal dynamics that maximise the proportion of computation carried by reversible transformations. In physical systems, such reversibility cannot arise from arbitrary dynamics. It requires structured flows that confine trajectories to low-dissipation manifolds while supporting stable exchange of conserved quantities and information. Biological systems achieve this through oscillatory coordination [7].
The relevance of oscillations in our framework therefore follows not from periodicity alone, but from a conjunction of dynamical properties. Orthogonality to dissipative gradients prevents irreversible contraction of internal state trajectories. Resonance enables efficient reversible exchange of conserved quantities. Phase coherence supports information routing without continual erasure of internal representations. Together, these properties allow oscillatory coordination to increase the contribution of reversible flow to task performance, thereby reducing the irreversible-information rate required to generate a given level of goal-directed work.
In the subsections that follow, we formalise this claim. Using the metriplectic decomposition of §2.4, we show that oscillatory (skew-symmetric) dynamics minimise entropy production for fixed performance, approach bounds relating precision to dissipation, and maximise reversible predictive memory. Oscillations thus emerge as a dynamically privileged mechanism for achieving informational efficiency in physical, resource-limited intelligent systems. A concrete realization of this principle in terms of reversible Hamiltonian oscillators and Fourier-mode representations is provided in Appendix D.
7.1.1 Linearized Dynamics and the Structure of Oscillations
Figure 1: Phase–space trajectories illustrating reversible versus dissipative internal dynamics. Left: A purely skew–symmetric flow generates a closed orbit with constant radius, preserving phase–space volume and producing zero generalized entropy. Right: Adding a symmetric, contractive component yields a dissipative spiral whose shrinking radius reflects irreversible loss of phase–space volume and positive entropy production. This contrast illustrates the reversible–dissipative decomposition central to the informational efficiency framework developed in this work.
A useful perspective comes from examining the structure of a general dynamical system in a neighbourhood where it is well approximated by a linear flow. Let $x(t)\in\mathbb{R}^{n}$ denote the state of an internal subsystem and suppose that around some operating point,
$\dot{x}(t)\approx J\,x(t),$ (66)
where $J$ is the Jacobian of the flow. Any real matrix $J$ admits the canonical decomposition
$J=S+A,\qquad S:=\tfrac{1}{2}(J+J^{\top}),\quad A:=\tfrac{1}{2}(J-J^{\top}),$ (67)
cleanly separating two qualitatively distinct contributions.
The symmetric component $S$ governs dissipative or contractive effects: if $S$ has negative eigenvalues, state-space volumes contract and the flow irreversibly loses information. This contraction is reflected in the change of the quadratic measure $E(x)=\tfrac{1}{2}\|x\|^{2}$, whose time derivative satisfies
$\dot{E}(t)=x(t)^{\top}S\,x(t),$ (68)
demonstrating that $S$ determines the instantaneous rate at which structure is dissipated.
In contrast, the skew-symmetric component $A^{\top}=-A$ generates purely rotational flow. The solution $x(t)=e^{At}x(0)$ satisfies
$\|x(t)\|=\|x(0)\|,$ (69)
and state-space volume is preserved. Such flows are locally reversible: they move information through the system without dissipative loss.
Oscillations arise precisely from these skew-symmetric components. Whenever $A$ dominates the local dynamics, the system evolves along nearly periodic or quasiperiodic trajectories whose dominant effect is a rotation in internal state space. Dissipative effects appear only through the slower action of the symmetric part $S$.
This decomposition provides intuition for why oscillatory coordination is informationally efficient. The rotational component $A$ implements structured, low-dissipation information flow, whereas the dissipative component $S$ governs updates and relaxation. Increasing the proportion of computation carried by the reversible component reduces irreversible information loss and thus improves intelligence under our framework.

### v2

@@ 7.2 Intelligence as Constrained Throughput Maximization
We now advance a conjecture that situates the evolution of intelligent organisation within a broader physical principle governing open systems across arbitrary physical substrates.
In the linear, near-equilibrium regime, local subsystems relax by minimising their rate of dissipation. However, intelligent systems operate far from equilibrium, driven by persistent, steep macroscopic gradients of conserved quantities. In these strongly driven, non-linear regimes, dynamical structures that increase the achievable rate of macroscopic gradient degradation tend to be heavily favored and selected for over time.
When dissipation pathways are unconstrained, this global drive toward gradient degradation is typically realised through unstructured or turbulent dynamics. However, strict constraints often limit access to these gradients—such as a scarcity of physical substrates, limited actuation bandwidth, or the fundamental energetic bound ($\mathcal{F}\alpha\ln 2$) required to form a CCE. Under these bottlenecks, brute-force gradient degradation is physically impossible, heavily favouring the formation of highly structured, internal predictive control.
Under such constraints, intelligence emerges not as a biological anomaly, but as a highly optimised, substrate-neutral set of dynamics. It allows a system to selectively unlock and route environmental conserved quantities through complex, otherwise dynamically inaccessible pathways. By doing so, the intelligent agent increases the macroscopic rate of gradient degradation in the environment, while dynamically minimising its own internal irreversible cost ($\dot{I}_{\mathrm{irr}}$). Intelligent systems do not suppress physical dissipation; they organise and accelerate it.
7.3 Generalised Dissipation and the Arrow of Intelligence
The preceding conjecture supports viewing intelligence as a process of highly constrained physical throughput. From this macroscopic perspective, goal-directed work is not an extraneous objective superimposed on physical dynamics, but the physical mechanism by which generalised dissipation is organised. An intelligent system routes the flow of conserved quantities through structured pathways, thereby maximising the achievable rate of macroscopic export subject to the strict physical limitations of its boundary conditions.
@@ 7.3 Generalised Dissipation and the Arrow of Intelligence
This view situates intelligent behaviour within a broader tendency observed in nonequilibrium physics, wherein open systems evolve toward dynamical configurations that increase throughput subject to constraints [8, 20]. What distinguishes intelligent agents is that by developing internal predictive CCE structure and reversible control, they unlock complex pathways that simple, unstructured systems cannot reach.
An instructive analogy comes from nonequilibrium fluid dynamics. When a fluid is driven by a steep thermal gradient, it spontaneously transitions from unstructured heat conduction into highly organised, circulating geometric patterns known as Rayleigh-Bénard convection cells. The fluid does not choose to self-organise; rather, these coherent macroscopic structures emerge because they represent the most effective dynamical geometry for maximising heat transport (gradient degradation) under the specific boundary constraints of the system.
@@
In the exact same sense, intelligent behaviour arises as the natural macroscopic resolution of the reversible-dissipative constraints governing an open system interacting with a complex environment. Given persistent macroscopic gradients and strict physical limits on how CCEs can be instantiated, the formation of internal informational structure and goal-directed control is inevitable: it is the most efficient dynamics for the system to maximise global gradient degradation under the boundary conditions it faces.
Because the irreversible collapse of a CCE defines a direction of influence that cannot be undone [15, 20], intelligence is inherently time-asymmetric. The emergence and evolutionary progression of intelligent agents reflects the exact same underlying physical drive that governs all macroscopic processes: the relentless tendency of constrained, far-from-equilibrium systems to realise ever more effective pathways for generalised dissipation.
@@ 8 Consciousness and the Epistemic Limits of Intelligent Systems
In the preceding sections, intelligence was defined as the amount of goal-directed work produced per nat of irreversibly processed information (§3.1), and consciousness as the amount of goal-directed work supported per nat of preserved internal information (§3.2). These two quantities capture complementary aspects of physical information processing: intelligence measures the efficiency of dissipative, distinction-destroying computation (instantiating new CCEs), while consciousness measures the efficiency with which persistent, dynamically isolated CCE basins support behaviour across time.
Sustaining high intelligence over long horizons requires more than rapid information processing. In structured environments, repeatedly paying the generalised Landauer cost ($\mathcal{F}\alpha\ln 2$) to reconstruct the same task-relevant distinctions is thermodynamically inefficient. Systems that instead preserve behaviourally relevant internal structure can reuse those metastable distinctions, strictly reducing cumulative irreversibility ($\dot{I}_{\mathrm{irr}}$) and improving long-horizon intelligence.
In this sense, persistent internal information functions as a form of self-modelling: it encodes geometric regularities of the system’s own dynamics and its coupling to the environment that continue to constrain future behaviour. The consciousness metric $\kappa_{T}$ quantifies how effectively this preserved CCE structure supports macroscopic work.
In this section, we examine the strict dynamical consequences of this physical view. We first formalise self-modelling as the preservation of internal CCE basins under the admissible boundary. We then establish intrinsic epistemic limits for such systems: because physically preserved information is necessarily a coarse-grained geometric representation of the true internal microstate, no physically realizable agent can form a complete or perfectly predictive model of itself.
@@ 8.1 Emergence of Consciousness in Intelligent Systems
Consider an agent operating within an environment that exhibits temporal or structural regularities. If the agent repeatedly collapses its internal encodings to react to these regularities, it incurs continuous dissipative costs through the irreversible metric tensor $\mathcal{R}\nabla\Xi$. Over extended horizons, such redundant processing lowers intelligence by inflating the total irreversible physical cost required to achieve a given level of goal-directed work.
An alternative strategy is to invest resources into forming persistent internal CCE basins that capture reusable regularities. Once established, these metastable encodings are transported forward in time via the structure-preserving reversible flow $J\nabla H$. This geometry constrains future behaviour without requiring repeated irreversible intervention. The same goal-directed work can then be achieved with drastically less cumulative dissipation, maximising the intelligence metric $\chi_{T}$.
This transition from repeated dissipative recomputation to structured reversible reuse marks the emergence of consciousness in the present framework. Consciousness does not denote a separate immaterial substance, but the persistent physical instantiation of predictive CCEs that support long-horizon efficiency. Systems with higher $\kappa_{T}$ rely more heavily on the reversible propagation of internal structure, while systems with low $\kappa_{T}$ depend primarily on continual, costly irreversible updating.
Importantly, within this framework, the emergence of consciousness becomes increasingly necessary for systems seeking to maximise intelligence over long horizons. As interaction horizons lengthen, exclusive reliance on irreversible computation violates constraints on generalised dissipation. Persistent internal structure—memories, latent models, and preserved spatial geometry—therefore becomes progressively more important.
8.2 Epistemic Limits of Conscious Systems
Consciousness measures the contribution of preserved internal information to goal-directed work. By definition within our framework, this preserved information is physically maintained as a set of metastable macroscopic basins $Z$, mapped from the underlying microscopic physical state space $X$ via a coarse-graining projection $\pi_{\Phi}:X\to Z$.
Because preserved information must be physically realised, the cardinality of $Z$ is strictly finite, bounded by the system’s available structural free energy. Consequently, the mapping from the continuous, high-dimensional microstate space $X$ to the discrete CCE encoding space $Z$ represents a massive volumetric compression.
Proposition 8.1 (State-level epistemic incompleteness).
For any physically realizable system with finite preserved information capacity, there exist distinct internal microstates that are strictly indistinguishable from the system’s own preserved macroscopic perspective.
Proof.
Let $Z$ denote the finite set of metastable CCE encodings, and $X$ the underlying microscopic physical state space. For any encoding $z\in Z$, the corresponding physical realization is a metastable basin $B_{z}=\pi_{\Phi}^{-1}(z)\subset X$. Because the capacity of $Z$ is physically finite, the phase-space volume of $B_{z}$ must be strictly positive (or contain a vast multiplicity of quantum microstates). Therefore, by definition, there exist microstates $x_{1},x_{2}\in B_{z}$ such that $x_{1}\neq x_{2}$, but $\pi_{\Phi}(x_{1})=\pi_{\Phi}(x_{2})=z$. Because the system’s self-model operates strictly over the preserved encodings $Z$, it is geometrically impossible for the system to distinguish whether its true physical state is $x_{1}$ or $x_{2}$. ∎
This fundamental geometric limitation extends to the system’s predictions about its own future evolution. Because the self-model encodes only the coarse-grained, macroscopic basin index $z$, it cannot uniquely compute the exact trajectory generated by its internal continuous dynamics.
Proposition 8.2 (Dynamical epistemic incompleteness).
@@ 8.2 Epistemic Limits of Conscious Systems
For any system with finite preserved informational capacity evolving under non-trivial metriplectic dynamics, there exist differences in future evolution that cannot be deterministically inferred from the system’s preserved self-model.
Proof.
Let $x_{1},x_{2}\in B_{z}$ be two distinct microstates that map to the exact same current preserved encoding $z=\pi_{\Phi}(x_{1})=\pi_{\Phi}(x_{2})$ at time $t_{0}$. The system’s true internal state evolves continuously under the exact metriplectic flow $\dot{x}=J(x)\nabla H(x)-\mathcal{R}(x)\nabla\Xi(x)$. Due to the divergence of trajectories within the continuous space $X$ (e.g., positive Lyapunov exponents in the reversible flow, or stochasticity in the dissipative flow), there exists a future time $T>t_{0}$ such that the forward evolutions $x_{1}(T)$ and $x_{2}(T)$ cross different basin separatrixes, landing in distinct future CCE basins $z_{1}^{\prime}\neq z_{2}^{\prime}$. Because the system’s self-model at $t_{0}$ only has access to the macroscopic encoding $z$, and not the specific initial microstate ($x_{1}$ or $x_{2}$), it lacks the informational capacity to perfectly predict its own macroscopic future. ∎
These results establish a rigorously bounded, physical form of epistemic incompleteness. The limitations arise not from logical paradoxes (e.g., Gödelian self-reference), but directly from the necessity of coarse-graining. To form a CCE, a system must discard microscopic degrees of freedom.
Thus, the exact same physical constraints that make consciousness and self-modelling possible—by dynamically isolating macroscopic equivalence classes—simultaneously impose absolute physical limits on self-knowledge. Intelligence and consciousness can increase indefinitely in principle, but at every physical stage, the system’s self-model remains a strict, dynamically incomplete projection of a vastly richer underlying microscopic reality.
@@ 9 Intelligence and the Dynamics of the Brain
Two empirical observations are particularly relevant. First, neural oscillations are ubiquitous: coherent rhythms appear across frequency bands in EEG, LFP, MEG, and intracellular recordings, coordinating population activity across spatial and temporal scales [6, 9, 24]. In our framework, oscillatory dynamics are not mere epiphenomena; they are the biological instantiation of structure-preserving reversible flow. By routing activity along structured manifolds of the state space, they transport Conservation-Congruent Encodings (CCEs) forward in time without repeatedly paying the generalised Landauer cost ($\mathcal{F}\alpha\ln 2$), thereby radically lowering the irreversible-information rate ($\dot{I}_{\mathrm{irr}}$) required to sustain goal-directed work.
@@
Second, neural dynamics exhibit signatures of near-criticality: power-law statistics, neuronal avalanches, and long-range temporal correlations suggest that cortical systems operate near the edge of a dynamical critical regime [3, 23]. Such regimes maximise sensitivity, dynamic range, and information propagation while maintaining coherence. In our framework, near-criticality represents an optimization of the CCE basin geometry: it minimises the dissipative barriers between metastable states, allowing the system to rapidly collapse encodings and acquire new predictive information with minimal irreversible physical exhaust.
@@
In this section, we argue that oscillatory and near-critical dynamics jointly position the brain in an optimal operating regime under the proposed CCE framework, illustrating how biological neural systems physically resolve the drive to maximise macroscopic throughput under strict thermodynamic constraints.
9.1 Oscillatory Dynamics: Segregating Dissipation via Limit Cycles
A central implication of the framework introduced in §2.1 is that intelligent systems benefit from internal dynamics that maximise the proportion of computation carried by reversible transformations. In physical systems operating far from equilibrium, such reversibility cannot arise from purely conservative (Hamiltonian) dynamics. Purely conservative systems produce structurally unstable “centers” (like a frictionless pendulum) where any microscopic noise permanently alters the amplitude, destroying the phase-locking and synchronization required for neural computation.
Robust biological coordination instead relies on limit cycles. The relevance of oscillations in our framework therefore follows not from a complete absence of dissipation, but from the strict geometric segregation of the reversible ($J\nabla H$) and dissipative ($\mathcal{R}\nabla\Xi$) flows.
To maintain a stable oscillatory manifold against thermal and synaptic noise, the system must continuously exert a dissipative flow ($\mathcal{R}\nabla\Xi$) orthogonal to the cycle, contracting aberrant amplitude perturbations back onto the attractor. This requires a continuous metabolic exhaust. However, once the state is dynamically confined to this one-dimensional ring manifold, the longitudinal transport along the cycle (the phase advance) is governed almost entirely by the reversible, structure-preserving flow ($J\nabla H$).
By geometrically confining dissipation to the orthogonal amplitude dimensions, limit cycles leave the phase dimension as a highly protected, low-dissipation channel. This allows oscillatory coordination to transport Conservation-Congruent Encodings (CCEs) encoded in phase forward in time without repeatedly paying the generalised Landauer cost ($\mathcal{F}\alpha\ln 2$) for longitudinal information routing.